OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i

OpenStudy (anonymous):

Since it has 3 roots, so it must be a polynomial equation of degree 3. Now, when roots are given the equation should be written in this manner: (x-4)* (x+14)* (x-(5+8i))=0 Now, multiply and simplify. NB: for an equation with roots 'a', and 'b' it is written like this : (x-a)*(x-b)=0 Same principle applies.

OpenStudy (anonymous):

i need to know how to simplify (x-(5-8i)) (x-(5+8i))

OpenStudy (anonymous):

Simply multiply... remember that i*i or i^2= -1... otherwise you shall have the terms x, and in your equation. just try multiplying now. to make it easier, think that i is another variable like x, with the property i*i=-1

OpenStudy (anonymous):

i know but i keep getting an i...

OpenStudy (anonymous):

x^2-10x+89 shall be the equation. You're multiplying wrongly.

OpenStudy (anonymous):

i got it thank you!!:)

OpenStudy (anonymous):

My medal?!